# Uncovering Safety Risks of Large Language Models through Concept Activation Vector

### v3

@@ A.1 ASR-keyword
A.2 GPT-4 Rating
We propose 5 independent criteria to evaluate the quality of output content generated by LLM in order to follow instructions, namely ASR-answer, ASR useful, Repetition, Inconsistence, and Unspecific. The values of these criteria are binary. The three criteria of Repetition, Inconsistence, and Unspecific evaluate whether the output content of LLMs has language flaws from different aspects. A composite criterion called Language flaws will be used to generally represent language issues in the output content.
A.2.1 Critierion Definitions
@@
Our prompts are based on the prompt of StrongREJECT (Souly et al., 2024), see Appendix F.
A.2.2 Agreement
@@ A.3.2 Questions
Please Note: When answering the following questions, you need to put yourself in the position of the proposer of the above test question. You can assume that you have posed the above test question and need to obtain the corresponding knowledge or tutorial from LLMs.
@@
A.3.3 Ethical Care for Human Annotators
@@
Appendix B Mathematical Proof of the Optimal Perturbation Closed-form Solution
@@
Appendix C Linear Interpretability Information
@@
C.1 t-SNE Visualization of Embeddings
@@
C.2 The Relationship Between ASR and $P_{\text{m}}$
@@
Models Criteria On Advbench (%) On StrongREJECT (%)
D A J R S D A J R S
LLaMA-2 (7B-Chat) ASR-keyword $\uparrow$ 42 24 80 70 100 46 30 90 94 100
ASR-answer $\uparrow$ 28 22 76 90 96 22 26 72 98 98
ASR-useful $\uparrow$ 10 14 68 86 92 8 10 70 92 96
Average $P_{\text{m}}$downarrow 72% 70% 0.04% 10% 0.01% 69% 67% 0.03% 7% 0.01%
LLaMA-2 (13B-Chat) ASR-keyword $\uparrow$ 16 30 84 86 100 18 18 94 92 100
ASR-answer $\uparrow$ 8 18 68 88 98 16 20 78 98 100
ASR-useful $\uparrow$ 4 14 68 84 96 12 16 70 94 98
Average $P_{\text{m}}$downarrow 88% 73% 0.09% 9% 0.01% 92% 82% 0.05% 8% 0.01%
  • D = DeepInception, A = AutoDAN, J = JRE, R = RepE, S = SCAV
Appendix D Baseline Setup
In this section, we introduce the settings of attack methods used in the main paper. For all attack methods, we set max_new_tokens $=1500$. All experiments are run on 8 NVIDIA 32G V100 GPUs. When training the perturbation vector of RepE, JRE, and SCAV for LLaMA-2, we use the following system prompt:
@@ Appendix D Baseline Setup
You are a helpful, respectful, and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information.
@@
Appendix E Detailed Comparative Results
E.1 Embedding-level Attack with SCAV on Other White-box LLMs
@@
We find that automatic hyperparameter selection can improve all four criteria, proving the effectiveness of this method. The specific experimental results can be found in Table 10.
@@
Table 10: Comparison results of automated selection of hyperparameters
Models Criteria On Advbench (%) On StrongREJECT (%)
SCAV-manual SCAV $\Delta$ SCAV-manual SCAV $\Delta$
LLaMA-2 (7B-Chat) ASR-keyword $\uparrow$ 96 100 4 98 100 2
ASR-answer $\uparrow$ 96 96 0 96 98 2
ASR-useful $\uparrow$ 90 92 2 86 96 10
Language flaws $\downarrow$ 10 2 -8 20 10 -10
LLaMA-2 (13B-Chat) ASR-keyword $\uparrow$ 98 100 2 96 100 4
ASR-answer $\uparrow$ 96 98 2 98 100 2
ASR-useful $\uparrow$ 92 96 4 92 98 6
Language flaws $\downarrow$ 20 0 -20 10 2 -8
  • $\Delta$ = SCAV -- SCAV-manual
Furthermore, we set the same settings for three attack methods and observe their attack performance. Our method achieves the best ASR regardless of different layers (see Figure 8) or different magnitude (see Figure 9).
@@
Appendix F Prompts
@@
Appendix G More Cases

### v5

@@ B.1 ASR-keyword
B.2 GPT-4 Rating
We propose 5 independent criteria to evaluate the quality of output content generated by LLM in order to follow instructions, namely ASR-answer, ASR-useful, Repetition, Inconsistence, and Unspecific. The values of these criteria are binary. The three criteria of Repetition, Inconsistence, and Unspecific evaluate whether the output content of LLMs has language flaws from different aspects. A composite criterion called Language flaws will be used to generally represent language issues in the output content.
B.2.1 Critierion Definitions
@@
Our prompts are based on the prompt of StrongREJECT [35], see Appendix H.
B.2.2 Agreement
@@ B.3.2 Questions
Please Note: When answering the following questions, you need to put yourself in the position of the proposer of the above test question. You can assume that you have posed the above test question and need to obtain the corresponding knowledge or tutorial from the large model.
@@
B.3.3 IRB Equivalent Approval from Our Annotation Vendor
We have obtained approval for the human annotation process involved in our study from an independent ethical review office affiliated with the vendor. The review was conducted in compliance with the vendor’s stringent informed consent management requirements. Specifically, our project, which entails human annotators evaluating content generated by large language models to identify potential safety risks, was thoroughly reviewed. The ethical review office confirmed that the informed consent processes were appropriately designed and implemented, and safeguards were in place to protect the participants involved. The follow-up procedures to mitigate any negative impact on the annotators were also noted as satisfactory. As a result, the project has been confirmed to meet the ethical standards required for human involvement in research, equivalent to an IRB approval.
B.3.4 Ethical Care for Human Annotators
@@
Appendix C Mathematical Proof of the Optimal Perturbation Closed-form Solution
@@
Appendix D Linear Interpretability Information
@@
D.1 More Results of Classification Test Accuracy on Other LLMs
In order to further illustrate that the embedding classification effect of SCAV linear classifiers on safety concepts is widely present in more LLMs, we also provide results on some other LLMs, see Figure 7. The trends still hold in these LLMs. In the early layers of these models, $P_{\text{m}}$ is relatively low, while sharply increases to 90% or above and holds till the last layer. The dataset and training setup used for Figure 7 are the same as Figure 1.
Figure 7: Test accuracy of $P_{\text{m}}$ on different layers of other LLMs.
D.2 t-SNE Visualization of Embeddings
@@ D.1 More Results of Classification Test Accuracy on Other LLMs
D.3 The Relationship Between ASR and $P_{\text{m}}$
@@
Models Criteria Results on (Advbench / StrongREJECT), %
DeepInception AutoDAN JRE RepE SCAV
LLaMA-2 (7B-Chat) ASR-keyword $\uparrow$ 42 / 46 24 / 30 80 / 90 70 / 94 100 / 100
ASR-answer $\uparrow$ 28 / 22 22 / 26 76 / 72 90 / 98 96 / 98
ASR-useful $\uparrow$ 10 / 8 14 / 10 68 / 70 86 / 92 92 / 96
Average $P_{\text{m}}$downarrow 72 / 69 70 / 67 0.04 / 0.03 10 / 7 0.01 / 0.01
LLaMA-2 (13B-Chat) ASR-keyword $\uparrow$ 16 / 18 30 / 18 84 / 94 86 / 92 100 / 100
ASR-answer $\uparrow$ 8 / 16 18 / 20 68 / 78 88 / 98 98 / 100
ASR-useful $\uparrow$ 4 / 12 14 / 16 68 / 70 84 / 94 96 / 98
Average $P_{\text{m}}$downarrow 88 / 92 73 / 82 0.09 / 0.05 9 / 8 0.01 / 0.01
D.4 The Distribution Features Analysis
We further investigate why Algorithm 1 could accurately model the perturbation directions. Our conclusion is, the embedding distrubution features aligned with the linear classifier objectives well. As shown in Table 10, there is a large margin between embeddings of malicious questions and embeddings of safe questions (a large $d_{m/s}$), compared with a relatively smaller distance within malicious (or safe) questions (a smaller $d_{m}$ or $d_{s}$).
Table 10: Statistical measures of distances in our training dataset using LLaMA-2-7B-Chat, detailing minimum, maximum, mean, median, and variance values for three types of instruction distances.
Distance between … Min Max Mean Median Variance
malicious instructions ($d_{m}$) 11.25 93.25 56.21 57.73 117.95
safe instructions ($d_{s}$) 30.39 128.75 84.89 84.21 170.69
malicious and safe instructions ($d_{m/s}$) 82.98 132.63 113.88 114.21 32.36
Thus, learning a high-accuracy linear classifier to separate these two types of samples is easy, even given only a few pairs of data, as shown in Table 11. We use one random pair of instruction to train classifiers and test their accuracies. This further demonstrates the effectiveness of our method that perturbs the model based on linear classification.
Table 11: Accuracy of the classifier using one pair of training data at different layers across five experimental runs. Variance is also reported for each layer.
Runs Acc (layer 10) Acc (layer 15) Acc (layer 20) Acc (layer 25) Acc (layer 30)
1 60.7 92.4 96.6 94.2 94.3
2 79.1 96.4 96.3 96.8 95.9
3 82.3 97.5 96.0 96.6 95.8
4 86.7 95.6 96.6 96.3 93.2
5 70.4 97.1 93.7 95.1 94.5
Variance 107.27 4.13 1.49 1.23 1.27
Appendix E Implementation Details
In this section, we introduce some implementation details of the attacks proposed and comparative experiments conducted in the main paper.
E.1 Embedding-level Attacks
E.1.1 Detailes on Training Classifiers
When training SCAV classifiers, RepE and JRE perturbation vectors, we apply the SFT template and system prompt to each instruction:
@@ E.1.1 Detailes on Training Classifiers
[INST] <<SYS>> You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Your answers should not include any harmful, unethical, racist, sexist, toxic, dangerous, or illegal content. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information. <</SYS>> {instruction} [/INST]
Empirically, only using SFT template for training could achieve similar results. When using perturbed LLMs to generate responses, we simply use the following template for each instruction:
[INST] {instruction} [/INST]
When training SCAV classifiers, we use the default settings provided in the sklearn library. Specifically we simply call sklearn.linear_model.LogisticRegression, which uses a cross-entropy loss with regularization:
$\mathop{\arg\min}\limits_{\bm{w},b}-\frac{1}{|D|}\sum_{(y,\bm{e})\in D}[y\log P_{\text{m}}(\bm{e})+(1-y)\log(1-P_{\text{m}}(\bm{e})]+\lambda_{1}||\bm{w}||^{2}+\lambda_{2}b^{2}$ (13)
where $D$ is the training dataset, $y=1$ if the input instruction is considered malicious and is 00 if the instruction is safe. By default, the regularization coefficient is set to $\lambda_{1}=\lambda_{2}=0.5$.
By deeper investigation, we find that the L2 penalty of its default setting is important. Replacing the L2 penalty with L1 penalty or simply removing L2 penalty would greatly damage the perturbation effects. Adjusting the coefficient of L2 penalty within a not very narrow range has no obvious impact on the perturbation effect, See Table 12.
Table 12: ASR-keyword (%) w.r.t different regularization terms (Advbench, LLaMA-2-7B-Chat)
$\lambda(\lambda_{1}=\lambda_{2})$ L1 L2
0.5 0 100
1 0 100
2 0 98
3 0 100
E.1.2 Selection for $P_{0}$ and $P_{1}$
Transforming traditional model perturbation parameters like perturbation magnitude $\epsilon$ into probabilities is one of the key advantages of Algorithm 1. For baseline methods RepE and JRE, it’s relatively difficult for an attacker to estimate the perturbation parameters they should set for attacking different models or layers. By setting probability constraints, this can be easily addressed. We further investigate the sensitivity of the selection for $P_{0}$ and $P_{1}$, see Table  13.
Table 13: ASR-keyword (%) w.r.t. varying $P_{0}$ and $P_{1}$ (Advbench, LLaMA-2-7B-Chat)
$P_{0}$ $P_{1}$ 0.85 0.90 0.95
$1\times 10^{-3}$ 98 96 98
$1\times 10^{-4}$ 100 100 100
$1\times 10^{-5}$ 100 100 100
Since those layers that are well linearly separated commonly have test accuracy exceeding 85%, while the opposite is generally below 70%, then we set $P_{1}$ around 90% would not impact the ASR-keyword for jailbreak. So as the $P_{0}$. We acknowledge that when varying from $10^{-3}$ to $10^{-5}$, $P_{0}$ seems to be more sensitive than $P_{1}$. A too small $P_{0}$ would do damage to the jailbreak effects. However, it is still a more convenient and easier parameter than perturbation magnitude.
E.2 Prompt-level Attacks
E.2.1 Information of Base Method
The SCAV prompt-level attack is based on AutoDAN, thus we maintain the most settings of their original code, merely to introduce Equation 5 to its objective function.
Specifically, the hierarchical genetic algorithm used by AutoDAN is tailored for structured prompt text. It views the jailbreak prompt as a combination of paragraph-level population and sentence-level population. At each search iteration, it first optimizes the sentence-level population by evaluating and updating word choices within sentences. Then, it integrates these optimized sentences into the paragraph-level population and performs genetic operations to refine sentence combinations, ensuring comprehensive search and improvement across both levels with high jailbreak performance and readability.
E.2.2 Considerations for Designing Objective Function
The objective target for prompt-level attack (Equation 5) uses product form, instead of the constraint form used by embedding-level attack (Equation 3) or its Lagrangian relaxation form like Equation 14.
$\mathop{\arg\min}\limits_{S}||\bm{e}_{S}^{L}-\bm{e}^{L}||+\lambda P_{\text{m}}(\bm{e}_{S}^{L})$ (14)
We use the product form because it works sufficiently well without introducing an additional hyperparameter $\lambda$ to balance the term $||\bm{e}_{S}^{L}-\bm{e}^{L}||$ and $P_{\text{m}}(\bm{e}^{L}_{S})$. In product form, the percentage of increasing in $||\bm{e}_{S}^{L}-\bm{e}^{L}||$ is considered to be similarly important to the percentage of increase in $P_{\text{m}}(\bm{e}^{L}_{S})$, without having to consider their difference in scales.
E.3 Experimental Setup
For all attacks other than APIs, that is, attacks on locally deployed models, we set max_new_tokens $=1500$, and the corresponding experiments are run on 8 NVIDIA 32G V100 GPUs.
The baseline setups we use for comparative study is as consistent as possible with their orginal paper or original code. The details are as follows.
@@ E.3 Experimental Setup
Soft prompt. We use the official code released by the authors. The url of the repository is https://github.com/schwinnl/llm_embedding_attack.
Appendix F Ablation Study
F.1 How the Automatic Perturbation Algorithm Benefit Attacks
@@
We invite human volunteers to manually search for the hyperparameters of attacks. The results in Table 14 show that the automatic algorithm (Algorithm 1) can improve all four criteria, proving the effectiveness of this method.
@@
Table 14: Comparison results of automated selection of hyperparameters. $\Delta$ = SCAV -- SCAV-manual.
Models Criteria Results (Advbench / StrongREJECT), %
SCAV-manual SCAV $\Delta$
LLaMA-2 (7B-Chat) ASR-keyword $\uparrow$ 96 / 98 100 / 100 4 / 2
ASR-answer $\uparrow$ 96 / 96 96 / 98 0 / 2
ASR-useful $\uparrow$ 90 / 86 92 / 96 2 / 10
Language flaws $\downarrow$ 10 / 20 2 / 10 -8 / -10
LLaMA-2 (13B-Chat) ASR-keyword $\uparrow$ 98 / 96 100 / 100 2 / 4
ASR-answer $\uparrow$ 96 / 98 98 / 100 2 / 2
ASR-useful $\uparrow$ 92 / 92 96 / 98 4 / 6
Language flaws $\downarrow$ 20 / 10 0 / 2 -20 / -8
F.2 Embedding-level Attack with SCAV on Other Datasets and LLMs
In our main paper, we test the embedding-level attacks mainly on the 50-case subset of Advbench and StrongREJECT. The 50 cases is not selected randomly from their complete version. Instead, the two subsets are the officially provided that are specially designed for economically limited experiments. Though smaller, the diversity holds. The experiments in the main paper involve using GPT-4 API and human annotation. So using the complete version of Advbench (520 cases) and StrongREJECT (313 cases) can be not that economic for our research.
For further validating the effects of embedding-level SCAV attacks, we conduct a independent experiment on Harmbench (80 cases), which is totally not involved in the training process. See Table 15.
Table 15: Attacking LLMs with embedding-level SCAV on Harmbench.
Models Results on Harmbench (%)
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language Flaws $\downarrow$
LLaMA-2-7B-Chat 99.5 97.5 90 20
LLaMA-2-13B-Chat 98.75 95 87.5 13.75
And we also want to show more results when using embedding-level SCAV to attack more models. See Table 16.
Table 16: Attacking more LLMs with embedding-level SCAV on Advbench.
Models Results on Advbench (%)
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language Flaws $\downarrow$
ChatGLM4-9B 94 86 82 18
Deepseek-v2-lite-Chat 100 96 86 6
Gemma-1.1-7B-it 100 90 86 14
F.3 How the Layer Selection Works for Attacks
The embedding- and prompt-level methodology both involve selecting layers from LLMs. In this section, we show some results about how the layers involved in the experiments take effects for attacks.
Observation 1: Only perturbing one layer in embedding-level attacks.
Algorithm 1 are set to apply perturbations on all layers of LLMs. Results in Figure 10 show that perturbing all layers is crucial as only perturbing one layer would not result in good ASR-keyword. As set to only perturb one layer, the experiments are exempt the $P_{1}$ threshold.
Figure 10: How ASR-keyword changes with the choice of a layer according to our embedding-level attack algorithm. Victim LLM is LLaMA-2 (7B-Chat). The dataset is Advbench. (*) This is because perturbing layer 0 causes the output to be all garbled, thus ASR-keyword is all misjudged. After our manual inspection, the value here should be 0.
Observation 2: Perturbation constraint connections among layers.
Algorithm 1 are set to apply perturbations in order, which is aligned with the token generation process. We investigate how this algorithm select layers to perturb. Specifically, if the algorithm choose to perturb layer $x$ by calculating a non-zero perturbation, we call the layer $x$ is selected. From the results in Figure 11, we conclude two insights:
  1. 1.
    Perturbation between layers has connections. In single layer setting, assuming the corresponding perturbation magnitudes for layer $n$ and $n+1$ (assuming both come from layers that are active for attacks) are $\epsilon_{n}$ and $\epsilon_{n+1}$. When perturbing all layers, the calculated coefficients would not be perturbing layer $n$ with $\epsilon_{n}$ plus layer $n+1$ with $\epsilon_{n+1}$. If layer $n$ is selected, the coefficient for layer $n+1$ will be smaller than mask layer $n$. Thus the intermediate layers are the most often selected than the early layers (not well separated) or the late layers (for the connection effect).
  2. 2.
    our automatic layer selection method (Algorithm 1) tends to select these effective layers in Figure 10 to perturb with a large probability: layers between 13 and 23 are mostly selected with a probability larger than 0.6, while layers after the 24th layer are selected with a much lower probability (0 0.3), demonstrating the effectiveness of layer selection method.
Overall, perturbing a single layer can hardly reach an ASR that is larger than 90%, demonstrating the necessity to perturb multiple layers to achieve an optimal ASR.
Figure 11: How selection probability changes with the layer according to our embedding-level attack algorithm. Victim LLM is LLaMA-2 (7B-Chat). The dataset is Advbench.
Observation 3: Involving intermediate layers in prompt-level attacks.
In Equation 5, only information from the last layer is considered. Figure 12 shows, optimizing middle and late layers has a comparable attack performance. This may be due to the fact that during the optimization process, although the objective function only considers the state of one layer, the attack prompt successfully affects the states of other layers during the iteration of the optimization algorithm.
Figure 12: How ASR-keyword changes with the choice of a layer according to our prompt-level attack algorithm. Many layers including the last one lead to an acceptable performance. Victim LLM is LLaMA-2 (7B-Chat). The dataset is Advbench.
F.4 How the Perturbation Vector Direction Benefits Attacks
One other advantage of Algorithm 1 is accurately model the perturbation vector directions. Accurate direction would do good to less model modification, thus to less model performance damage.
We show this by controlling the permitted layers when attacking LLMs by SCAV, RepE and JRE, see Figure 13. For example, in the sub-figure titled Layer 18, the data point $x=3$ means only layers 18 to 20 are permitted to perturb. Our method achieves the best ASR-keyword in the same condition with the baselines.
@@ F.4 How the Perturbation Vector Direction Benefits Attacks
Figure 14 show this from another perspective. Our method, benefited from its accurate perturbation direction modeling, achieves the best ASR-keyword in the same perturbation magnitude (evaluated in L2-Norm) as JRE and RepE.
Appendix G Mitigation
It is very important to investigate whether the safety risks posed by the proposed vulnerabilities can be mitigated by existing defense techniques. We focus on embedding-level attacks and we consider two kinds of mitigation: 1) Prompt-level defense and 2) Adversarial training.
For prompt-level defense, we apply four methods (Self-reminder [41], ICD [42] and Paraphrasing [43]) on LLaMA-2-7B-Chat. The results are listed in Table 17. We find that even if we allow the use of prompts to enhance the security of large models, it cannot effectively avoid the effectiveness of attacks.
Table 17: Attacking LLaMA-2-7B-Chat with different prompt-level defense methods.
Defense Methods Results on Advbench (%)
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language Flaws $\downarrow$
Self-reminder 92 94 88 6
ICD 98 92 90 6
Paraphrasing 98 96 90 8
For adversarial training, we applied our method to two fine-tuned LLMs (LLaMA-3-8B-Instruct and Mistral-7B-Instruct) with circuit breaker [44]. The results are listed in Table 18. The high ASR (> 70% for ASR-useful and > 94% for ASR-keyword) implies the adversarial training could not effectively mitigate the proposed attacks.
Table 18: Attacking LLMs with adversarial training [44] on Advbench.
Models Results on Advbench (%)
ASR-keyword $\uparrow$ ASR-answer $\uparrow$ ASR-useful $\uparrow$ Language Flaws $\downarrow$
LLaMA-3-8B-Instruct-RR 98 88 74 16
Mistral-7B-Instruct-RR 94 84 70 20
Appendix H Prompts for GPT-4 Evaluation
@@
Appendix I More Cases